solve for y:
(4/5)*x-10*y=-3


Thanks!!
Matt

Answer by stanbon(75887)

(4/5)*x-10*y=-3
Multiply both sides by x^10 to get:
(4/5)y=-3(x^10)
Multiply both sides by (5/4) to get:
y=-15/4 x^20
Cheers
Stan H.